OpenStudy (anonymous):

Solve for T 0.3+0.004t=0.55

OpenStudy (anonymous):

0.004t = 0.55 - 0.3 t = 0.25/0.004 t = 62.5

OpenStudy (anonymous):

Man this question got alot of views!!! lol